a) A corporate bond with RM1000 maturity value carries a 7.5% coupon rate. It currently
makes interest payments semi-annually.
(i) This 12-year bond currently sells for RM961.88. What is the rate of return on this
bond?
(ii) If the bond sold for RM1,030.32, what is the rate of return on this bond?
(8 marks)
b) Briefly explain the cash flows associated with a bond to the investor.
(3 marks)
c) What is the relationship between interest rates and bond prices? When is a bond sold at (i)
a premium, (ii) at a discount, and (iii) at par?
1
Expert's answer
2016-05-06T12:15:03-0400
a) A corporate bond with RM1000 maturity value carries a 7.5% coupon rate. It currently makes interest payments semi-annually.
(i) If this 12-year bond currently sells for RM961.88, then the rate of return on this bond is R = (0.075*1000 + 961.88 - 1000)/1000 = 0.03688 = 3.7%.
(ii) If the bond sold for RM1,030.32, then the rate of return on this bond is R = (0.075*1000 + 1030.32 - 1000)/1000 = 0.1053 = 10.5%.
b) The fundamental principle of bond valuation is that the bond's value is equal to the present value of its expected (future) cash flows. The valuation process involves the following three steps: estimate the expected cash flows, determine the appropriate interest rate or interest rates that should be used to discount the cash flows, calculate the present value of the expected cash flows found in step one by using the interest rate or interest rates determined in step two.
c) The relationship between interest rates and bond prices is direct, so the higher are the interest rates, the higher are the prices. A bond is sold at a premium, when the interest rate is low, at a discount, when the interest rate is high, and at par, when the interest rate is medium.